OpenStudy (anonymous):

How do you get x out of the exponent?? 3^5x = 27^ (4+2x)

OpenStudy (anonymous):

rewrite \[27=3^3\] so \[27^{4+2x}=(3^3)^{4+2x}=3^{12+6x}\]

OpenStudy (anonymous):

then ignore the exponential and put \[5x=12+6x\] and solve for x